OpenStudy (anonymous):

@iGreen @AnswerMyQuestions Maybe 3 or 4 more then im done. Thanks for the help so far.

OpenStudy (anonymous):

A cone-shaped piece of baker’s chocolate has a radius of 3 cm and a height of 9 cm. Samia shaves off some of the chocolate. The remaining piece is still shaped like a cone with a radius of 3 cm, but the height is 6 cm. How much chocolate did Samia shave off? Use 3.14 to approximate pi and express your final answer in hundredths. _____ cm^3

OpenStudy (igreen):

Find the volume of the two cones and subtract.

OpenStudy (igreen):

\(V = \dfrac{1}{3} \pi r^2 h\) Plug in 3 for the radius and 9 for the height and 3.14 for pi.

OpenStudy (anonymous):

1/3 x 3.14 x 3^2 x 9 = 84.78

OpenStudy (igreen):

Yep, now for the other one: \(V = \dfrac{1}{3} \pi r^2 h\) Radius = 3 Height = 6

OpenStudy (anonymous):

56.52

OpenStudy (igreen):

Yep, now subtract 84.78 - 56.52

OpenStudy (anonymous):

28.26

OpenStudy (igreen):

Yep, that's our answer.

OpenStudy (anonymous):

Thanks :)